### v2

# CP-Agent: Agentic Constraint Programming

Abstract
Translating natural language into formal constraint models requires expertise in the problem domain and modeling frameworks. To investigate whether constraint modeling benefits from agentic workflows, we introduce CP-Agent, a Python coding agent using the ReAct framework with a persistent IPython kernel. Domain knowledge is provided through a project prompt of under 50 lines. The agent iteratively executes code, observes the solver’s feedback, and refines models based on the execution results.
We evaluate CP-Agent on CP-Bench’s 101 constraint programming problems. We clarified the benchmark to address systematic ambiguities in problem specifications and errors in ground-truth models. On the clarified benchmark, CP-Agent solves all 101 problems. Ablation studies indicate that minimal guidance outperforms detailed procedural scaffolding, and that explicit task management tools have mixed effects on focused modeling tasks.
1 Introduction
Constraint Programming (CP) is a declarative paradigm for solving combinatorial problems where users specify what constraints a solution must satisfy, rather than prescribing how to find it [26]. The primary limitation preventing wider adoption of CP remains the modeling bottleneck: translating natural language problem descriptions into formal constraint models requires deep expertise in both the problem domain and the modeling framework [5]. This challenge has been characterized as the “holy grail” of constraint programming [24].
Large Language Models (LLMs) are now being applied to automate this translation process. The CP-Bench benchmark [14], introduced to evaluate LLM-driven constraint modeling capabilities, contains 101 diverse combinatorial problems sourced from various established CP community resources. These problems encompass optimization and satisfaction tasks, representing a variety of distinct constraint types. This benchmark addresses limitations of existing datasets, such as NL4Opt [17] and Logic Grid Puzzles [25], which lack the diversity and complexity necessary for evaluating constraint modeling capabilities.
The CP-Bench study [14] evaluated multiple LLMs across three modeling frameworks. With direct prompting, the best models achieved up to 65% accuracy using Python-based frameworks (CPMpy, OR-Tools), while MiniZinc (a popular constraint modeling language) peaked around 50%. Applying inference-time methods (repeated sampling combined with self-verification) to CPMpy improved accuracy to approximately 70%.
The choice between fixed and agentic workflows is well-established in the literature. As Wei et al. [29] demonstrated with chain-of-thought prompting, and subsequent work by Yao et al. [32] and Schick et al. [19] confirmed that one should use a fixed, single-shot workflow for well-scoped, short-horizon tasks, and prefer an agentic, tool-using loop as task complexity, multi-step reasoning, or external information needs increase. We hypothesize that constraint modeling from natural language falls into the latter category, requiring an agentic workflow with iterative refinement, hypothesis testing, and dynamic adaptation based on solver feedback. Our results confirm this hypothesis.
To test this hypothesis, we introduce CP-Agent, a constraint modeling agent based on our general-purpose agentic-python-coder system. CP-Agent operates within a persistent IPython kernel [16], enabling stateful and iterative development. Rather than embedding domain-specific logic into its architecture, the agent is guided by a concise project prompt. This design allows the agent to execute code, observe runtime feedback from the constraint solver, and adjust models based on execution results.
The underlying coding agent is also available as an MCP server [1]. The Model Context Protocol (MCP), now governed by the Linux Foundation, has become a de facto standard for connecting language models with external tools [8]. This integration enables interactive constraint modeling within MCP-compatible clients such as Claude Desktop or Cursor, where users can guide the agent through complex modeling decisions.
On the CP-Bench dataset of 101 problems (with added clarifications for some ambiguous problems and output format requirements), CP-Agent solves all problems correctly. While our primary reference language model is Claude Sonnet 4.5, we also tested this approach with several other models from four different providers. Swapping only the project prompt enables the agent to adapt to new domains, as confirmed by follow-up work [23] on logic programming modeling tasks.
We provide an experimental analysis of an explicit task management tool within our agentic coding workflow. The results are mixed: while task tracking helped on some previously failing problems, it also broke problems that worked without it. The net effect was negative (6 fewer problems solved) with significantly increased token usage, suggesting that for focused, single-file generation tasks, the overhead can distract as often as it organizes.
2 Related Work
CP-Agent aligns with the emerging neurosymbolic paradigm where neural language understanding is combined with symbolic constraint satisfaction [18]. The LLM-based agent translates natural language into formal constraints, while the CP solver provides a sound and complete search over the solution space.
Recent work has explored LLM agents for software engineering tasks. SWE-agent [31] resolves GitHub issues through a file-system-based Agent-Computer Interface with commands for navigating and editing repository files. CP-Agent focuses instead on single-script generation using a stateful IPython kernel, eliminating file I/O overhead for this class of problems. CodeAct [28] argues for using executable Python code as a unified action space for LLM agents, showing that this approach outperforms more restrictive formats. Our work applies this principle to constraint modeling, demonstrating that a general action space can solve specialized tasks by utilizing focused prompts. The iterative refinement in CP-Agent relates to automated program repair work. ChatRepair [30] introduced conversational repair, where an LLM iteratively refines patches based on test failures. RepairAgent [3] treats repair as an autonomous task with tools for code search, patching, and testing. Our agent performs similar iterative refinement but for generative tasks, translating natural language to constraint programs, rather than repairing existing code.
Benchmark quality affects evaluation validity. EvalPlus [13] demonstrated that popular benchmarks contain insufficient tests, with augmented test suites revealing previously undetected errors. ClassEval [4] addressed structural complexity through manual benchmark curation. Our work on CP-Bench addresses problem specification ambiguity, enabling more rigorous measurement of constraint modeling abilities.
3 Agentic Python Coder
CP-Agent is based on a generic Python coding agent equipped with a special prompt for constraint modeling. The coding agent implements a minimal yet capable coding agent using the ReAct framework [32]. The system architecture separates domain-agnostic agent capabilities from task-specific knowledge, enabling rapid adaptation to new problem domains through prompt engineering rather than code modification. The implementation of the generic coding agent is available from a GitHub repository and via PiPy for installation via “pip install agentic-python-coder” [20].
3.1 Architecture
Virtual execution space
Unlike software engineering agents that navigate complex file trees, this architecture targets single-script development, where all state is contained in a single file. The agent operates in a virtual execution space provided by the IPython kernel [16], which maintains a persistent state across multiple code executions. Variables, functions, and imports remain accessible throughout a session, eliminating file I/O overhead while enabling rapid hypothesis testing. The agent operates interactively via a Read–Eval–Print-Loop (REPL), and benefits from immediate feedback to all code manipulations. This design is well-suited for constraint programming models and similar focused generation tasks.
To manage this persistent state and guide the execution flow, we employ LangGraph [12]. This framework manages the agent’s workflow, enabling the development of stateful agents through durable execution and memory management. The framework supports the ReAct pattern [32] of interleaved reasoning and acting, allowing the agent to interact with the execution environment and adjust solutions through iterative cycles. The agent consists of three main components: the ReAct loop controller, which operates the reasoning-action cycle; a minimal tool suite focused on code generation and execution; and a project prompt that provides domain-specific guidance. The ReAct loop (Figure 1) manages the iterative cycle of reasoning about the task, selecting and executing tools, and observing results until the agent produces a final solution.
TaskReasonAct (Tools)Solutionexecutecompleteobserve
Figure 1: The ReAct (Reason and Act) framework employed by the agent.
Tool suite
The agent operates with a minimal, focused tool set. The core tool is python_exec(code: str), which executes Python code in the persistent kernel environment and returns stdout/stderr for solver feedback. Each execution builds on the previous state, allowing the agent to develop solutions incrementally. Calling save_code signals completion and writes the final solution to disk; a maximum iteration limit prevents runaway execution.
Kernel implementation
The system launches an IPython kernel as a separate subprocess, providing process isolation while enabling state persistence. All Python state resides in this long-running kernel, creating a stateful coding environment similar to an interactive session where each execution builds upon previous ones.
Dynamic dependency injection
Rather than having the agent execute pip install statements for adding required Python libraries or relying on pre-built container images, the system implements a custom kernel manager that wraps the standard IPython kernel launch. Using the uv package manager [2], the system injects requested packages directly into an ephemeral environment at kernel startup. This approach provides process-level isolation: each session runs in a clean environment with only the specified dependencies, while the kernel maintains state persistence across executions. The design decouples the agent’s reasoning from dependency mechanics, thereby eliminating the latency associated with runtime package installation.
Task management
The agent can use a task management tool to create and maintain an in-memory list of tasks, commonly referred to as a “todo list.” The tool call uses the format todo_write(todos: List[Dict[str, Any]]). Each task must contain id, content, status, and priority fields, with only one task marked as “in_progress” at any time.
LLM integration
The agent connects to language models through OpenRouter, [^1] https://openrouter.ai providing access without direct API dependencies. This approach allows switching between different models by changing a single configuration parameter. The system streams responses with usage statistics and detailed tracking of tool calls. These logs provide insights into the agent’s problem-solving process and resource consumption.
Prompt architecture
The agent operates with a three-layer prompt hierarchy. The system prompt (approximately 100 lines) defines general coding agent behavior and tool usage patterns without task-specific specialization, instructing the agent on how to reason about code, execute iterative development cycles, and utilize available tools effectively. The project prompt provides domain-specific knowledge (such as constraint modeling), examples, and best practices relevant to a particular problem class. The task prompt contains the specific problem to solve, typically read from a task.md file in the working directory or provided directly via a command line parameter. This separation enables the same agent infrastructure to tackle diverse problem domains by swapping project prompts while maintaining consistent operational behavior.
4 CP-Agent
For constraint programming, the generic coding agent is equipped with a domain-specific project prompt and uses CPMpy (see below) as the modeling framework. The approach relies on the agent’s ability to use general coding capabilities guided by domain expertise expressed in the project prompt, rather than specialized constraint programming logic embedded in the architecture. The agent can test simplified problem instances, verify constraint satisfaction on example data, implement custom validation logic, and refine models based on solver feedback. This flexibility, combined with the persistent execution environment, enables the agent to discover and correct modeling errors that would escape detection in single-pass generation approaches.
4.1 CPMpy
The agent uses CPMpy [7] as the constraint programming foundation. This Python-based library builds on NumPy for efficient handling of decision variables as arrays, allowing integration with the broader Python ecosystem. Its solver-agnostic design enables modeling problems once and solving them using various backends, including CP, MIP, SMT, and SAT solvers. The choice of Python matches findings from Michailidis et al. [14] that LLMs perform better with Python-based frameworks than domain-specific languages, likely due to the billions of lines of Python code in their training data.
4.2 Project Prompt
The project prompt (cpmpy.md) is under 50 lines and provides focused guidance for constraint modeling. The complete prompt is included in the GitHub repository [20].
The prompt contains five core rules: use CPMpy’s declarative modeling rather than implementing search algorithms, output only valid JSON using json.dumps(), always import the json module when outputting JSON, check the exact output format required, and manually test solutions to ensure they satisfy the problem requirements. A basic code template demonstrates the standard structure, and brief sections cover three essential CPMpy constraints (AllDifferent, sum, Circuit) and optimization via model.minimize() and model.maximize().
This minimal prompt resulted from an ablation study. Initial experiments utilized a substantially longer prompt, approximately 800 lines in length, which contained detailed modeling procedures, extensive problem-type catalogs, and debugging guidance. The shorter version achieved equal or better performance, suggesting that for modern LLMs with constraint programming knowledge in their training data, concise guidance focusing on common pitfalls suffices—excessive detail may limit problem-solving flexibility.
5 Benchmark Problems
The CP-Bench dataset [14] provides a benchmark for evaluating automated constraint modeling capabilities. It contains 101 constraint programming problems drawn from four established sources: CSPLib [6], CPMpy examples [7], models by Kjellerstrand [11], and problems from the APLAI course at KU Leuven. The problems span diverse combinatorial classes, including scheduling, assignment, sequencing, graph problems, and combinatorial design, with 30 optimization problems and 71 decision problems providing balanced coverage of constraint programming challenges.
Each problem is presented as a natural language description in Markdown format, along with input data and required output variables. Solutions must execute within a specified time limit and produce results in JSON format, using the specified variable names. CP-Bench includes ground-truth CPMpy models for all problems, enabling semantic validation through constraint satisfaction checking rather than relying solely on syntactic correctness or execution success. This semantic validation is essential for rigorous evaluation of automated modeling approaches. The ground-truth models are withheld from the agent and used solely for post-hoc verification of generated solutions.
5.1 Fixed Workflow Results
Michailidis et al. [14] evaluated fixed workflow approaches on CP-Bench, testing progressively detailed prompts (basic instructions, modeling guidelines, API documentation) and retrieval-augmented in-context learning. The highest accuracy of 70% was achieved by combining repeated sampling with self-verification using GPT-4-turbo. These fixed workflow approaches share a common limitation: they cannot test hypotheses, debug errors, or verify partial solutions.
5.2 Benchmark Corrections and Clarifications
Our initial evaluation attempts revealed issues with the original benchmark that prevented rigorous automated evaluation. The CP-Bench authors independently acknowledged222personal communication, August 2025 several problems, including ground-truth models with symmetry-breaking constraints that reject valid solutions, unclear output format requirements, and ambiguities in problem statements. Without clear specifications, evaluation conflates the ability to model constraint satisfaction problems with the ability to guess intended formats and interpret ambiguous descriptions. We added clarifications to isolate the actual modeling capability. These clarifications are necessary to maintain focus on constraint modeling rather than resolving ambiguity or making guesses.
Our benchmark updates fall into three categories. First, we clarified 31 ambiguous problem statements. For instance, Problem 080 (TSP) required specifying whether Euclidean distances should be rounded or truncated—different methods lead to different distance matrices and thus different optimal solutions. For Problem 013 (magic square), we corrected a misleading description that stated the goal is to find a “magic square” when the problem actually requires a “magic sequence.” Problem 060 (crossword) provided a visual grid and word list, but required explicit specification of slot lengths and intersection constraints to be solvable. Second, we added structured JSON schemas to all problems to clarify the required output format. Third, we corrected errors in 19 ground-truth models: removing symmetry-breaking constraints from 10 and fixing logic errors in 9 others.
The modifications do not make the constraint modeling problems easier—they remove confounding factors that obscure actual modeling capability. The underlying combinatorial structure and search space complexity remained unchanged; no changes hint at constraint formulations or modeling strategies. The success rates reported in the original CP-Bench paper were obtained using the original benchmark and ground-truth models. Direct comparison between our results and theirs would be inappropriate, as they measure different capabilities. With clear specifications, other methods would likely show higher success rates as well. We document all changes in supplementary materials and provide the clarified benchmark for fair comparison in future work.
6 Experiment
6.1 Experimental Setup
Our experimental setup executed all 101 problems using CP-Agent, implemented as agentic-python-coder (version 2.2.1), with Claude Sonnet 4.5 (version 20250929). Each problem execution produces a solution file containing the generated CPMpy model and a conversation log recording the complete agent interaction. We validated the solutions by running them against the ground-truth CPMpy models. For satisfaction problems, we verified that the solutions satisfy all ground-truth constraints. For optimization problems, we also compared the objective value to confirm optimality. This approach correctly handles problems with multiple valid solutions while ensuring correctness and accuracy. We ran each experiment three times to account for stochastic variance.
The agentic-python-coder and all experimental data, including problem specifications, ground-truth models, generated solutions, and execution logs, are available online [22]. The data is encrypted (password Ckc1ADsKXRfmzD2) to prevent training data contamination.
6.2 Results
CP-Agent achieved 100% accuracy on all 101 CP-Bench problems, successfully solving all problems across all three independent runs. All generated solutions were validated against the reference CPMpy models and produced correct and optimal solutions.
Table 1: Tool usage, token consumption, and execution time for all 101 CP-Bench problems (averaged over three independent runs). Column ex shows the number of python_exec calls (iterative refinement cycles). Columns in/out show input/output tokens in thousands. Column time shows execution time in seconds (rounded).
Problem ex in out time Problem ex in out time
001_car_seq 2.7 24 2 43 052_circular_t 3.7 29 2 47
002_template 2 21 2 52 053_clock_trip 1 14 1 26
003_autocorr 3.3 26 2 44 054_cmo_2012 3 25 3 49
004_golomb 2.3 22 2 42 055_coin3_appl 2 18 2 34
005_all_inter 1.7 19 2 32 056_coins_grid 4.3 38 2 47
006_vessel 2.3 26 3 51 057_contractin 1 14 1 25
007_perfect_sq 4 30 2 44 058_covering_o 2 24 2 42
008_social_gol 7.7 54 3 61 059_crew 2.3 36 4 62
009_basketball 23.7 256 12 245 060_crossword 4 40 3 55
010_nonogram 8 90 6 110 061_crypta 4.7 67 4 74
011_progressive 8 79 5 93 062_eighteen_h 1 12 1 21
012_schurs 2.7 23 2 36 063_facility_l 3 29 3 57
013_magic_sq 1 15 1 22 064_fifty_puzz 1.3 15 1 23
014_crossfig 28.3 737 33 542 065_three_coin 1 15 1 27
015_langford 1.7 19 2 36 066_three_sum 1 13 1 20
016_sports 8.3 75 5 99 067_twelve_pac 1.3 14 1 25
017_bibd 2.7 24 2 37 068_bus_schedu 1 14 1 27
018_word_des 6.3 47 3 80 069_jobshop 3.3 29 3 52
019_steiner 3.7 29 2 41 070_knapsack 1 13 1 22
020_num_part 2.7 21 2 60 071_mario 7.3 86 6 110
021_diamond 4.3 41 4 77 072_minesweepe 5 37 3 57
022_graceful 1 16 1 26 073_n_puzzle 10.7 106 6 119
023_n_queens 2 20 2 42 074_packing_re 8.3 75 5 83
024_costas 2 20 1 30 075_resource_c 5.7 68 4 78
025_hadamard 4.3 31 2 46 076_room_assig 2 21 2 38
026_abbots_puz 1 13 1 23 077_send_more 1 14 1 25
027_added_corn 2 20 2 37 078_set_game 3.3 34 3 60
028_ages_of_th 4 27 2 53 079_sudoku 2 23 2 39
029_age_changi 4.7 38 3 57 080_tsp 3.3 24 2 40
030_allergy 2 21 2 44 081_who_killed 2 21 2 42
031_among 2 17 1 42 082_wolf_goat 3.3 28 3 58
032_appointmen 1 14 1 36 083_zebra 3.7 38 4 59
033_archery_pu 2.3 18 1 49 084_bank_card 1 13 1 23
034_arch_frien 2 24 3 56 085_climbing_s 1.3 16 1 31
035_assignment 2.7 22 2 49 086_color_simp 2 18 1 30
036_autoref 4.3 31 2 46 087_exodus 5 45 4 71
037_bales_of_h 4 33 3 66 088_farmer_and 2.3 19 2 34
038_bananas 1.7 18 2 39 089_five_floor 1 14 1 24
039_best_host 1.7 18 2 33 090_grocery 4.3 31 2 51
040_big_bang2 1.7 21 2 39 091_guards_and 2 20 2 37
041_bin_packin 1 13 1 23 092_hardy_1729 1.7 15 1 28
042_birthday_c 1 14 1 26 093_kidney_exc 3.3 27 2 44
043_bowls_and 1.7 16 1 30 094_magic_sq 1.3 15 1 28
044_broken_wei 2 19 2 40 095_maximal_in 1.7 19 2 36
045_building_b 2.3 20 2 34 096_money_chan 1 13 1 24
046_cabling 1.3 17 2 36 097_movie_sche 2.7 23 2 41
047_calvin_puz 2.7 24 2 49 098_people_in 1 15 2 36
048_candies 1.3 16 2 28 099_subsets_10 1 14 1 27
049_capital_bu 1 14 1 25 100_subset_sum 2.7 21 1 37
050_chess_set 6.3 79 3 59 101_thick_as_t 2.7 26 3 54
051_circling_s 2.3 25 3 47
Token consumption averaged 37k input and 3k output tokens per problem. Problem 014 (crossfig) represents an extreme outlier, consuming 737k input tokens due to its complex crossword-figure structure, which requires extensive iterative constraint refinement.
Execution times ranged from 20 seconds to 542 seconds, with an average of 52 seconds per problem across the three runs. The timing variation between the three independent runs (averaging 48, 57, and 52 seconds, respectively) demonstrates the stochastic nature of LLM-based systems.
Table 1 summarizes computational cost; solution quality is perfect across all runs. The agent called python_exec between 1 and 28 times per problem, with simpler problems requiring fewer iterations and complex scheduling or optimization problems requiring more exploration. Problem 014 (crossfig) required the most Python executions (28 average), followed by 009 (basketball) with 24 executions, reflecting their combinatorial complexity.
To illustrate how the agent approaches constraint modeling, we examine two representative runs that demonstrate different problem-solving patterns.
6.3 Problem 061: Cryptarithmetic Puzzle
The cryptarithmetic problem required assigning unique digits (0–9) to letters in a 20-digit addition equation, ensuring the arithmetic holds while each letter maps to a distinct digit. The agent’s solution process demonstrates strategic adaptation when an initial approach proves unworkable.
The first four executions built the foundation: defining ten integer variables (one per letter), adding an AllDifferent constraint, mapping variables to their positions in the three numbers, and constraining leading digits to be non-zero. This setup proceeded smoothly, establishing the basic problem structure.
The fifth execution attempted to model the problem as a single unit by treating each 20-digit number as a single large integer variable. The agent constructed expressions like num1 = letter[0]*10**19 + letter[1]*10**18 + ... for all three numbers and added a constraint num1 + num2 == result. When the model tried to solve, the solver raised a TypeError: __init__(): incompatible constructor arguments because variable domains with bounds near $10^{20}$ exceeded the solver’s integer representation capacity.
Recognizing the fundamental incompatibility between this modeling approach and solver limitations, the agent abandoned the large-integer strategy entirely. The agent redesigned the model, introducing binary carry variables and modeling addition column by column, thereby mimicking manual arithmetic:
carries = intvar(0, 1, shape=21)
for pos in range(20):
digit1 = letter_to_var[num1_chars[19-pos]]
digit2 = letter_to_var[num2_chars[19-pos]]
result_digit = letter_to_var[result_chars[19-pos]]
carry_in = carries[pos]
carry_out = carries[pos+1]
model += (digit1 + digit2 + carry_in
== result_digit + 10 * carry_out)
This column-wise formulation ensured that all intermediate values remained within standard integer bounds, enabling the solver to determine the correct digit assignment.
6.4 Problem 082: Wolf, Goat, and Cabbage
While the above example demonstrates error recovery, this river-crossing puzzle illustrates the agent’s behavior when proceeding without failures. The agent built the temporal model through four distinct phases.
The first phase established state variables: wolf_pos[t], goat_pos[t], cabbage_pos[t], and boat_pos[t], which track locations (shore 0 or 1) across an 8-step timeline. The second phase fixed boundary conditions—all entities start at shore 0 and end at shore 1. The third phase encoded the movement mechanics: the boat crosses at each step, the farmer transports at most one item per trip, and item movements coincide with boat movements. The fourth phase added safety constraints:
for i in range(steps):
model += ~((goat_pos[i] == cabbage_pos[i])
& (goat_pos[i] != boat_pos[i]))
model += ~((wolf_pos[i] == goat_pos[i])
& (wolf_pos[i] != boat_pos[i]))
These constraints forbid the goat and cabbage (or wolf and goat) from sharing a shore without the farmer present. The systematic layering—state, boundaries, movement, and safety—reflects a deliberate construction process where the agent tests each phase before proceeding to the next. The model solved correctly on the first execution.
These two problems reveal complementary behaviors: complete redesigns when facing fundamental obstacles (Cryptarithmetic) and methodical incremental construction when the path is clear (Wolf, Goat, and Cabbage). The agent adapts its approach rather than following a fixed procedure.
7 Model Generalization
To evaluate whether the approach generalizes beyond Claude Sonnet 4.5, we tested six frontier models on ten representative problems with three runs each. The models span four providers: Anthropic (Sonnet 4.5, Opus 4.5, Haiku 4.5), OpenAI (GPT-5.1-codex-mini), xAI (Grok-code-fast-1), and DeepSeek (v3.2). We excluded Google’s Gemini models due to infrastructure issues during the experimental period. All models used a temperature of 0.0 with a 20-minute timeout per problem.
Table 2 shows the results. All tested models achieved a success rate of 80–100% using this architecture. Sonnet 4.5 maintained perfect accuracy, while Opus 4.5 and GPT-5.1 each reached 97%. Haiku 4.5, Grok, and DeepSeek achieved 80–87% accuracy, with most failures occurring on Problem 009 (basketball), although sporadic failures also occurred on other problems. Basketball requires coordinating nine constraint categories, including mirroring schemes, home/away patterns, rival pairings, and date-specific requirements—the highest constraint density in this subset.
Some runs hit the 20-minute timeout before completion. Review of execution logs confirms these agents were not converging toward correct solutions, suggesting the timeout did not cut short otherwise successful attempts.
These results show the architecture generalizes across model families. Success correlates with model capability rather than provider: the agentic workflow provides value at all capability levels tested. Prompt optimization frameworks, such as DSPy [10], could automatically tune the project prompt for specific models, potentially narrowing the performance gap observed with other models.
Table 2: Multi-model evaluation (3 runs each). Numbers indicate successful runs out of 3. Model success rates: Sonnet 100%, Opus/GPT-5 97%, Haiku 87%, DeepSeek 83%, Grok 80%.
Problem
Sonnet 4.5
Opus 4.5
Haiku 4.5
GPT-5.1
Grok
DeepSeek
001_car_seq 3 3 3 3 3 3
002_template 3 2 3 3 3 2
004_golomb 3 3 3 3 2 3
007_perfect_sq 3 3 3 3 3 3
009_basketball 3 3 1 2 0 0
010_nonogram 3 3 3 3 3 2
013_magic_sq 3 3 3 3 3 3
015_langford 3 3 1 3 2 3
020_num_part 3 3 3 3 3 3
037_bales_of_h 3 3 3 3 2 3
8 Task Management Tool
Explicit task tracking for LLM agents—maintaining a mutable list of pending, in-progress, and completed tasks—was introduced by BabyAGI [15] and formalized in plan-and-solve prompting [27]. The todo_write tool in our coder implements this pattern. Our architecture requires only todo_write, not a corresponding todo_read: without history compaction, all previous tool calls remain in the conversation context, so the agent can always see its earlier task updates. To evaluate whether explicit task management benefits constraint modeling, we conducted an ablation study comparing performance with and without this tool.
We used a smaller model (Claude Haiku 4.5) than our default model (Claude Sonnet 4.5) to reveal potential differences. We selected 40 problems, stratified by difficulty: 20 easy problems (solved in 27 seconds or less by Sonnet 4.5) and 20 hard problems (requiring 58 seconds or more). For each problem, we executed three independent runs with todo_write disabled and three with it enabled. Table 3 presents the results.
Table 3: Task management ablation on 40 problems (Claude Haiku 4.5, 3 runs per condition). Columns: p = success (out of 3), tok = tokens (K, averaged), t = time (s, averaged). Bold highlights problems where success changed.
no todo todo no todo todo
Easy Problem p tok t p tok t Hard Problem p tok t p tok t
013_magic_sq 3 15 26 3 39 35 008_social_gol 3 36 36 3 122 74
022_graceful 3 38 32 3 49 42 009_basketball 3 752 303 0 181 119
026_abbots_puz 3 35 29 3 57 40 010_nonogram 3 47 48 3 112 67
041_bin_packin 3 10 13 3 52 48 011_progressive 1 201 111 2 730 197
042_birthday_c 3 19 20 3 52 40 014_crossfig 0 1986 617 2 1592 492
049_capital_bu 3 10 18 3 37 41 016_sports 0 112 70 1 712 228
053_clock_trip 2 17 18 3 87 54 018_word_des 3 41 40 3 160 90
057_contractin 3 16 18 3 38 35 021_diamond 3 69 63 3 130 77
062_eighteen_h 3 15 16 3 57 49 037_bales_of_h 3 40 40 3 149 81
064_fifty_puzz 3 10 12 3 35 29 050_chess_set 3 10 16 3 38 31
065_three_coin 3 15 16 3 70 60 059_crew 3 80 55 0 137 75
066_three_sum 3 9 13 3 31 26 061_crypta 3 43 31 2 297 127
067_twelve_pac 3 9 12 3 39 33 071_mario 3 243 145 2 291 135
068_bus_schedu 3 18 20 3 50 39 073_n_puzzle 3 203 95 3 273 122
070_knapsack 3 9 11 3 37 32 074_packing_re 3 22 25 3 134 81
077_send_more 3 11 13 3 41 37 075_resource_c 3 55 51 3 86 61
084_bank_card 3 10 16 3 26 24 078_set_game 3 24 24 3 105 70
089_five_floor 3 11 14 3 33 30 082_wolf_goat 3 62 60 0 286 156
096_money_chan 3 13 22 3 32 34 083_zebra 3 30 32 3 150 91
099_subsets_10 3 16 19 3 28 26 087_exodus 3 123 88 3 223 106
Easy (60 runs) 59 60 Hard (60 runs) 52 45
On easy problems, todo_write increased token consumption by 191% (from 15K to 44K average) and execution time by 111% (from 18s to 38s), while improving success by just one run (59/60 to 60/60). On hard problems, todo_write reduced success from 52/60 to 45/60 while still adding overhead: 41% more tokens (209K to 295K) and 27% longer execution (98s to 124s).
Examining individual problems reveals the pattern. The task management tool helped three previously failing hard problems—crossfig, sports, and progressive—by providing structure for extended problem-solving. However, it broke three problems that worked without it: basketball, crew, and wolf_goat all dropped from 3/3 to 0/3. The task-tracking overhead appears to distract the agent from solutions it would otherwise find.
9 Discussion
Our experiments yield several key observations regarding the design and evaluation of agentic systems for constraint modeling.
The ablation study, which shows that a short prompt outperforms an 800-line version, suggests that for foundation models with constraint programming knowledge in their training data, the prompt’s role is to set high-level strategy and details on output format rather than provide exhaustive procedural instructions. Prompt compression research supports this finding: shorter, denser prompts can match or exceed verbose prompts [9]. Excessive detail may limit the agent’s ability to solve problems effectively.
The primary evaluation used Claude Sonnet 4.5. Multi-model testing on a representative 10-problem subset confirms that the architecture generalizes: six models from four providers achieved success rates of 80–100%, with performance correlating with model capability.
This pattern extends to built-in planning mechanisms. The optional todo_write tool, intended to help the agent manage tasks, reduced overall success when tested with a smaller model. The effect was asymmetric: todo_write helped on already-failing problems but broke problems that worked without it. For problems where the entire solution fits in a single script, the overhead of maintaining an explicit task list is more distracting than beneficial. When state is fully captured in code and execution output, the agent’s internal reasoning seems to suffice without externalized planning.
10 MCP Server
The Model Context Protocol (MCP) provides a standardized interface for connecting language models with external tools and data sources [8]. Introduced by Anthropic in late 2024 [1], MCP has become a de facto standard for agent-tool interaction, with adoption by major AI platforms including OpenAI and integration into development environments such as Cursor, Claude Code, VS Code (via Cline), Windsurf, and OpenAI Codex CLI. In December 2025, MCP governance transferred to the Agentic AI Foundation under the Linux Foundation.
While our standalone CP-Agent excels at batch processing, interactive development benefits from human guidance through complex modeling decisions. We expose the same persistent IPython kernel as an MCP server (ipython_mcp) for integration with Model Context Protocol-compatible clients such as Claude Desktop or Cursor. The server provides python_exec for code execution, the same tool central to our standalone agent, along with three tools for controlling and monitoring the coding infrastructure: python_reset to clear state and install packages to the virtual environment, python_status to query session information, and python_interrupt to halt long-running computations. This interactive mode also serves an educational purpose: users can observe how natural language constraints are formalized into CPMpy models, request explanations, and refine models based on solver feedback. Once developed, the generated code can be saved for deployment in other contexts.
Our MCP integration complements the system MCP-Solver [21], which provides specialized tools for constraint modeling (MiniZinc, PySAT, Z3, ASP). In contrast, the ipython_mcp server offers general-purpose Python execution, where domain adaptation occurs through prompts rather than tool specialization.
11 Conclusion
CP-Agent solves all 101 problems in the clarified CP-Bench benchmark, demonstrating that agentic workflows with execution feedback can reliably automate constraint modeling from natural language.
The agent architecture combines the ReAct framework with a persistent IPython kernel. Our ablation study shows that a minimal project prompt is sufficient, and that explicit task management tools generally add overhead for focused modeling tasks. However, they can help maintain focus during longer problem-solving sessions. For foundation models with relevant domain knowledge, concise guidance and stateful execution prove more effective than detailed procedural scaffolding or explicit planning mechanisms.
The architecture’s separation of domain expertise (prompts) from execution infrastructure (the agent) allows adaptation to other scientific computing domains. Follow-up work on ASP-Bench [23] demonstrated excellent performance on 128 answer set programming problems, utilizing the identical architecture with only a domain-specific project prompt, thereby confirming that our approach generalizes across declarative programming paradigms.
Acknowledgments
This research was supported by the Austrian Science Fund (FWF) projects 10.55776/P36420 and 10.55776/COE12.